# Separate and Conquer: Decoupling Co-occurrence via Decomposition and Representation for Weakly Supervised Semantic Segmentation

### 29 Feb 2024 v2

@@ 3.5 Representation with Patch Semantics
Local Semantics Extraction. Following the memory setup in contrastive learning, we store patch semantics across the dataset. However, simply taking two views of a patch as positive pair is not helpful to learn the difference among co-contexts. We propose a category tag pool to match the memory bank and guide the contrast among co-categories. Both patch semantics and category tags are stored as supportive knowledge to decouple the co-context at patch level. Specifically, we build a local teacher to extract features from history patches and update the reservoir and tag pool chronologically by storing the most recent key semantics ${k}$ and its corresponding tag $t_{i}$ while dequeuing the oldest. Mathematically, given the input $(x,t)$, the current query and key embeddings with tags are denoted as $B_{q}$ and $B_{k}$, the oldest are $B_{-q}$ and $B_{-k}$, respectively. Then the reservoir paired with tags is defined as:
@@
where $N_{v}=\sum_{k_{+}\in R(x,t_{i})_{+}}\mathbb{1}\left(q_{i}^{T}k_{+}<\mu\left(q_{i}\right)\right)$ is the number of noisy pairs.
@@ 3.6 Training Objectives
${L}_{{SeCo}}={L}_{{cls}}+{L}_{{cls}}^{{aux}}+\alpha{L}_{LiG}+\beta{L}_{LiL},$ (9)
It is noted that the proposed framework SeCo generates the pseudo masks online and is trained end-to-end to achieve the dense segmentation task. The loss for segmentation adopts cross-entropy loss ${L}_{{seg}}$. Thus, the overall loss is: ${L}={L}_{{SeCo}}+\gamma{L}_{{seg}}$. Following previous approaches [36, 37, 40], we leverage regularization losses to enforce the spatial consistency of CAM and the predicted masks.
@@
Figure 3: Qualitative segmentation results of AFA [36], ToCo [37] and ours on VOC and COCO. SeCo differentiate co-occurrence precisely.
@@ 4.1 Experimental Settings
Datasets and Evaluation Metrics. The proposed method is evaluated on PASCAL VOC 2012 [14] and MS COCO 2014 [26]. PASCAL VOC contains 21 semantic categories. Following the practice [29, 36, 37], we use the augmented dataset with $10,582$, $1,449$, and $1,456$ images for training, validating, and testing, respectively. MS COCO includes 81 classes. $82,081$ images are used for training, and $40,137$ images are used for validating. Mean Intersection-Over-Union (mIoU) is used as evaluation criteria. Confusion ratio, i.e., the number of false positive pixels / that of true positives, is designed to evaluate the efficacy on suppressing false positives from co-occurrence.
@@ 4.2 Main Results
Evaluation of CAM and Pseudo Masks. Fig. 2(b) (a) quantitatively reports the quality of the initial CAM and pseudo masks generated by SeCo and other recent competitors on VOC train set. It shows that SeCo generates better CAM seeds with $74.8\%$ mIou, even surpassing other methods refined with post-processing. With CRF post-refinement [18], the quality of the generated pseudo masks further improves to $76.5\%$, significantly higher than both single-staged and multi-staged methods by at least $4.3\%$.
Performance of Semantic Segmentation. Fig. 2(b) (a) reports the performance of the semantic segmentation of SeCo on VOC val and test set. The proposed SeCo achieves $74.0\%$ and $73.8\%$ mIoU on val set and test set, respectively. Fig. 2(b) (b) compares our segmentation performance with other recent methods on COCO val set. Without external data, SeCo achieves $46.7\%$ mIoU outperforms multi-staged  [27] competitors tackling co-occurrence and single-staged [37] SOTAs by $1.3\%$ and $4.4\%$ mIoU, respectively.
Prediction results on VOC and COCO are visualized in Fig. 3. It illustrates that SeCo can precisely localise the co-occurring objects on both datasets. For example, our method is capable of filtering the distracting backgrounds (water, railroad) from objects, or differentiating the co-occurring foregrounds (horse, person and bicycle), which demonstrates the competence at tackling co-occurrence.
@@ 5.1 Efficacy of Key Components
Ablative experiments on the key components of SeCo are conducted. Tab. 2 shows the segmentation results on VOC val set. Here, w/o LiG means no category knowledge is extracted, w/o LiL means that the local semantic reservoir and category tag pool are not maintained, and w/o Tag Rec. means the tag rectification is not incorporated. The category knowledge in LiG acts as class centroids for training while the patch knowledge makes the category semantics more compact, which facilitate the difference among co-contexts. As can be seen, without LiG in SeCo, the precision and recall drop heavily by $4.8$ and $3.9$, respectively. It verifies that the proposed method can effectively help to suppress the false positives and generate more complete masks. LiL in SeCo works the same. Without it, the precision, recall and mIoU drops heavily. The tag rectification improves precision from $82.6$ to $84.0$. It works by guaranteeing the right constructions of positive pairs and negative pairs, reducing the noise in the contrastive representation.
Figure 4: CAMs for co-contexts on VOC between SeCo and competitors [36, 37]. SeCo accurately activates targets (star).
@@ 5.2 Effectiveness of Tackling Co-occurrence
In Tab. 3, representative co-occurring pairs (e.g., {Boat, water}, {train, railroad}, etc.) in VOC val set are selected to validate the the effectiveness of our method to tackle co-occurrence. The metrics of IoU and confusion ratio (in bracket) are adopted. Confusion ratio is calculated by FP/TP, the lower the better. It reports that SeCo demonstrates significantly lower confusion ratio, such as boat ( $-79\%$ to ToCo), train ( $-21\%$), and higher IoU in all representative co-occurring pairs. For all categories, SeCo surpasses other competitors with $0.23$ confusion ratio, lower by at least $9\%$, which shows the superiority of our method to suppress the false positives from co-occurrence.
@@ 5.3 Further Analysis
Hyper-parameter Sensitive Analysis. We use grid search strategy to explore the hyper-parameters. The analysis of key parameters, such as patch size, loss weights, reservoir capacity, EMA momentum, temperature factors, etc., is specifically discussed in Supplementary Materials. SeCo remains consistent with variations of hyper-parameters, which shows the robustness of our method.
Training Efficiency Analysis. SeCo is designed in a single-staged paradigm to efficiently tackle the co-occurrence issue. The training efficiency comparisons are reported in Tab. 4. CLIMS [46] leverages CLIP model to tackle co-occurrence and consists of 3 steps. Since each step can not be implemented in parallel, it takes $1068$ minutes to finish the WSSS workflow. Compared to it, SeCo takes $417$ minutes to finish the workflow and outperforms it tackling co-occurrence by a significant margin. Notably, SeCo achieves more favorable performance compared to other single-staged competitors [36, 37] as well.
@@
Feature Representation Analysis. The feature representation is enhanced to decouple the co-occurrence. Here, we further visualize the co-context feature representation at image level to validate the effectiveness. As shown in the right of Fig. 5, we compute the similarity among the category knowledge from images. It is observed that each prototype in the knowledge is only highly related to itself, which suggests that the co-occurring semantics are separated contributed by our method. As shown in the left of Fig. 5, the visualization with t-SNE [41] also validates the efficacy.
In addition, we visualize the co-context feature representation at patch level as well. We decompose a demo image with co-occurring objects {dining table, chair} into $16$ patches, as shown in the left of Fig. 6. Each star (index from $1$ to $16$) denotes a patch, orange stars for table semantics and blue ones for chair semantics. We calculate the similarity between patches and category prototypes. As shown in the right of Fig. 6, the patch semantics has the correct relationship to the corresponding prototypes. Moreover, the patches with index $(3,7,10)$ containing both table and chair semantics shows a high relationship to both category information. This validates that SeCo can build deeply fine-grained representation to recognize co-contexts.
@@
Figure 6: Patch feature representation. Left: sample image with co-contexts {table, chair}. The orange star represents table patch and the blue is chair patch. Right: similarity among patch semantics and category prototypes. The x-axis means the patch index and the y-axis is the $20$ category prototypes on VOC.
@@ 6 Conclusion
One limitation is that, although we make the attempt to allocate tags and reduce the bias, it inevitably remains co-category patches and allocates wrong tags. In the future, leveraging patches with adaptive size or adopting other denoising techniques is a potential research for WSSS.

### 21 Mar 2024 v3

@@ 3.5 Representation with Patch Semantics
Local Semantics Extraction. Following the memory setup in contrastive learning, we store patch semantics across the dataset. However, simply taking two views of a patch as a positive pair is not helpful to learn the difference among co-contexts. We propose a category tag pool to match the memory bank and guide the contrast among co-categories. Both patch semantics and category tags are stored as supportive knowledge to decouple the co-context at patch level. Specifically, we build a local teacher to extract features from history patches and update the reservoir and tag pool chronologically by storing the most recent key semantics ${k}$ and its corresponding tag $t_{i}$ while dequeuing the oldest. Mathematically, given the input $(x,t)$, the current query and key embeddings with tags are denoted as $B_{q}$ and $B_{k}$, the oldest are $B_{-q}$ and $B_{-k}$, respectively. Then the reservoir paired with tags is defined as:
@@
where $N_{v}=\sum_{k_{+}\in R(x,t_{i})_{+}}\mathbb{1}\left(q_{i}^{T}k_{+}<\mu\left(q_{i},t_{i}\right)\right)$ is the number of noisy pairs.
@@ 3.6 Training Objectives
${L}_{{SeCo}}={L}_{{cls}}+{L}_{{cls}}^{{aux}}+\alpha{L}_{LiG}+\beta{L}_{LiL}.$ (9)
It is noted that the proposed framework SeCo generates the pseudo masks online and is trained end-to-end to achieve the dense segmentation task. The loss for segmentation adopts cross-entropy loss ${L}_{{seg}}$. Thus, the overall loss is: ${L}={L}_{{SeCo}}+\gamma{L}_{{seg}}$. Following previous approaches [36, 37, 40], we leverage regularization losses to enforce the spatial consistency of CAMs and the predicted masks.
@@
Figure 3: Qualitative segmentation results of AFA [36], ToCo [37] and ours on VOC and COCO. SeCo differentiates co-contexts precisely.
@@ 4.1 Experimental Settings
Datasets and Evaluation Metrics. The proposed method is evaluated on PASCAL VOC 2012 [14] and MS COCO 2014 [26]. PASCAL VOC contains 21 semantic categories. Following the practice [29, 36, 37], we use the augmented dataset with $10,582$, $1,449$, and $1,456$ images for training, validating, and testing, respectively. MS COCO includes 81 classes. $82,081$ images are used for training, and $40,137$ images are used for validating. Mean Intersection-Over-Union (mIoU) is used as evaluation criteria. Confusion ratio, i.e., the number of false positive pixels / that of true positives, is designed to evaluate the efficacy of suppressing false positives from co-occurrence.
@@ 4.2 Main Results
Evaluation of CAMs and Pseudo Masks. Fig. 2(b) (a) quantitatively reports the quality of the initial CAMs and pseudo masks generated by SeCo and other recent competitors on VOC train set. It shows that SeCo generates better CAM seeds with $74.8\%$ mIou, even surpassing other methods refined with post-processing [18]. With the simple multi-scale refinement [37], the quality of pseudo masks further improves to $76.5\%$, significantly higher than both single-staged and multi-staged methods by at least $4.3\%$.
Performance of Semantic Segmentation. Fig. 2(b) (a) reports the performance of the semantic segmentation of SeCo on VOC. The proposed SeCo achieves $74.0\%$ and $73.8\%$ mIoU on the val set and test set, respectively. Fig. 2(b) (b) compares our segmentation performance with other recent methods on COCO val set. Without external data, SeCo achieves $46.7\%$ mIoU and outperforms multi-staged  [27] competitors tackling co-occurrence and single-staged [37] SOTAs by $1.3\%$ and $4.4\%$ mIoU, respectively.
Prediction results on VOC and COCO are visualized in Fig. 3. It illustrates that SeCo can precisely localize the co-occurring objects on both datasets. For example, our method is capable of filtering the distracting backgrounds (water, railroad) from objects, or differentiating the co-occurring foregrounds (horse, person and bicycle), which demonstrates the competence at addressing co-occurrence.
@@ 5.1 Efficacy of Key Components
Ablative experiments on the key components of SeCo are conducted. Tab. 2 shows the segmentation results on VOC val set. Here, w/o LiG means no category prototypes is extracted, w/o LiL means that the local semantic reservoir and category tag pool are not maintained, and w/o Tag Rec. means the tag rectification is not incorporated. The category prototypes from G-teacher act as class centroids for training while the patch knowledge from L-teacher makes the category semantics more compact, which facilitates the difference among co-contexts. As can be seen, without LiG, the precision and recall drop heavily by $4.8\%$ and $3.9\%$, respectively. It verifies that the proposed method can effectively help suppress the false positives and generate more complete masks. Without LiL, the precision, recall and mIoU drop heavily as well. The tag rectification improves precision from $82.6\%$ to $84.0\%$. It works by guaranteeing the right constructions of positive pairs and negative pairs, reducing the noise in the contrastive representation.
Figure 4: CAMs for co-contexts on VOC between SeCo and competitors [36, 37]. SeCo accurately activates the targets (star).
@@ 5.2 Effectiveness of Tackling Co-occurrence
In Tab. 3, representative co-occurring pairs (e.g., {Boat, water}, {train, railroad}, etc.) in VOC val set are selected to validate the effectiveness of our method to tackle co-occurrence. The metrics of IoU and confusion ratio (in bracket) are adopted. Confusion ratio is calculated by FP/TP, the lower the better. It reports that SeCo demonstrates significantly lower confusion ratio than ToCo, such as boat ( $-79\%$), train ( $-21\%$), and higher IoU in all representative co-occurring pairs. For all categories, SeCo surpasses other competitors with $0.23$ confusion ratio, lower by at least $9\%$, which shows the superiority of our method to suppress the false positives from co-occurrence.
@@ 5.3 Further Analysis
Hyper-parameter Sensitive Analysis. The analysis of key parameters, such as patch size, loss weights, reservoir capacity, EMA momentum, temperature factors, etc., is specifically discussed in Supplementary Materials.
Training Efficiency Analysis. SeCo is designed in a single-staged paradigm to efficiently tackle the co-occurrence issue. The training efficiency comparisons are reported in Tab. 4. CLIMS [46] leverages CLIP model to tackle co-occurrence and consists of 3 progressive steps, which takes $1068$ minutes to finish the WSSS workflow. Compared to it, SeCo takes $417$ minutes to finish the workflow and outperforms it by a significant margin. Notably, SeCo achieves more favorable performance compared to other single-staged competitors [36, 37] as well.
@@
Feature Representation Analysis. We visualize the co-context feature representation at image level to validate the effectiveness. As shown in the right of Fig. 5, we compute the similarity among the category knowledge from images. It is observed that each prototype in the knowledge is only highly related to itself, which suggests that the co-occurring semantics is separated. As shown in the left of Fig. 5, the visualization with t-SNE [41] also validates the efficacy.
In addition, we further visualize the co-context feature representation at patch level. We decompose a demo image with co-occurring objects {dining table, chair} into $16$ patches, as shown in the left of Fig. 6. Each star (index from $1$ to $16$) denotes a patch, orange stars for table semantics and blue ones for chair semantics. We calculate the similarity between patches and category prototypes. As shown in the right of Fig. 6, the patch semantics has the correct relationship to the corresponding prototypes. Moreover, the patches with index $(3,7,10)$ containing both table and chair semantics show a high relationship to both category information. This validates that SeCo can successfully recognize co-contexts instead of being biased to the one or another.
@@
Figure 6: Patch feature representation. Left: sample image with co-contexts {table, chair}. The orange star represents the table patch and the blue is the chair patch. Right: similarity among patch semantics and category prototypes. The x-axis means the patch index and the y-axis is the $20$ category prototypes on VOC.
@@ 6 Conclusion
One limitation is that, although we make the attempt to allocate tags and reduce the bias, it inevitably remains co-category patches and allocates wrong tags. In the future, leveraging patches with adaptive size or adopting other denoising techniques is the potential research for WSSS.
7 Acknowledgement
This work was supported by National Natural Science Foundation of China (82072021), Shanghai Municipal Science and Technology Committee (23410710400, 22YF1409300).